# Memory-Dominated Quantum Criticality as a Universal Route to High-Temperature Superconductivity

### v1

@@ Abstract
Equation (54) reproduces the empirical Uemura scaling observed across cuprates and other unconventional superconductors. In contrast to BCS superconductors, where phase stiffness is large and irrelevant to $T_{c}$, correlated systems operate in a regime where infrared collective dynamics simultaneously controls pairing and coherence. In the present framework, this is not an additional assumption but follows directly from the single infrared control parameter $\rho_{0}$.
@@
Such long–time tails naturally generate 1/f1/f–type noise spectra and anomalously slow relaxation, phenomena widely reported in strange metals and cuprate superconductors. Unlike conventional overdamped criticality, where correlations decay exponentially and dynamics is effectively Markovian, the memory–dominated regime predicts non-Markovian response with persistent temporal structure. Because the same slow–mode reservoir controls both pairing enhancement and dynamical response, superconducting and normal–state anomalies become two manifestations of a single infrared spectral reorganization.
@@
The TDOS controls: (i) the extent of the slow–mode reservoir away from criticality [Eq. (47)], (ii) the dynamical amplification of intrinsic pairing tendencies [Eq. (50)], (iii) the superfluid stiffness and phase coherence [Eq. (54)], and (iv) the emergence of non-Markovian long–time response. Quantum criticality thus acts not merely as a source of fluctuations but as a self–organized dynamical amplifier that simultaneously governs thermodynamic and dynamical phenomena in strongly correlated superconductors.
This unified picture yields a concrete strategy for experimental tests: systems that exhibit enhanced low–frequency TDOS weight (or equivalently, strong long–time correlations) should simultaneously display increased pairing susceptibility, elevated $T_{c}$, and reduced phase stiffness consistent with Uemura scaling. The resulting real-frequency spectral function generated by the memory-dominated self-energy, including its universal scaling form and infrared structure, is derived in Appendix E.
@@
The central result of this work is that infrared collective dynamics in strongly correlated quantum matter is governed not by local Markovian dissipation but by the time–scale density of states of collective relaxation modes. Criticality corresponds to a reorganization of the entire dynamical spectrum, and universality classes are naturally classified by the low–energy scaling of the TDOS. In particular, a finite TDOS at vanishing relaxation rate defines a memory–dominated infrared fixed point characterized by long–time kernels $K(t)\sim 1/t$.
@@
This perspective clarifies the limited domain of validity of conventional coarse–grained field theories. Away from criticality, relaxation spectra are gapped in the infrared, leading to finite memory times and justifying reduction to local-in-time damping and a small number of collective order parameters. Near correlated critical regimes, however, relaxation rates condense into a continuous spectrum extending to zero. Internal modes cannot be integrated out: the full relaxation spectrum remains dynamically active at all time scales. The breakdown of Markovian coarse–graining is therefore a universal consequence of dynamical criticality rather than a model-dependent artifact.
This observation explains why Hertz–Millis–type approaches, which assume overdamped local dynamics of a small set of collective fields, cannot capture the infrared structure of strongly correlated critical matter. Criticality does not correspond to the softening of a single bosonic mode but to a spectral condensation of near–marginal collective degrees of freedom. The TDOS thus emerges as the natural organizing variable of infrared dynamics, replacing phenomenological damping parameters and providing a microscopic foundation for anomalous relaxation and transport.
@@
A central physical consequence of this spectral reorganization is the qualitative enhancement of superconducting instabilities. Rather than requiring finely tuned bosonic mediators, superconductivity emerges generically from dynamical amplification of intrinsic electronic pairing tendencies by the slow–mode reservoir. The exponential suppression characteristic of weak-coupling BCS theory is replaced by algebraic scaling controlled by infrared spectral weight.
@@
A single infrared parameter, $\rho_{0}$, governs thermodynamic and dynamical phenomena alike. The superconducting dome reflects truncation of long–time memory away from criticality, Uemura scaling follows from the shared spectral origin of pairing and phase coherence, and anomalous normal-state dynamics arises from the same non-Markovian kernel. Quantum criticality therefore acts as a self-organized dynamical amplifier that simultaneously controls fluctuations, transport, and superconducting order.
Although electrons propagate as Bloch quasiparticles on a crystalline lattice, collective particle–hole and particle–particle excitations form extended many-body coordinates not constrained to lattice energy eigenstates. In strongly correlated regimes these collective modes are generically overdamped by the electronic continuum, rendering a relaxation-rate description more fundamental than an energy-eigenmode picture. While conventional lattice theories organize collective behavior in terms of Hamiltonian eigenmodes, the present framework shows that near criticality the relevant organizing structure is instead the dynamically self-organized spectrum of relaxation rates.
This relaxation–rate organization is not a phenomenological assumption but a direct consequence of the fundamental structure of nonequilibrium quantum and classical dynamics. As shown in the Appendix F, both Liouvillian evolution in open quantum systems and generalized memory–kernel dynamics naturally decompose into continua of decay modes characterized by complex relaxation eigenvalues rather than by Hamiltonian energy eigenstates.
Even in the absence of spatial degrees of freedom, the Laplace–domain representation of non-Markovian dynamics generically yields continuous spectra of relaxation rates whenever long–time memory kernels are present. The emergence of a dense TDOS therefore reflects a universal mathematical property of dynamical systems with extended temporal correlations, not a model-dependent artifact of coarse–graining.
@@
From this perspective, energy eigenmodes describe conservative reversible dynamics, while relaxation spectra encode the physically relevant irreversible organization governing collective behavior near criticality. Strongly correlated quantum matter near dynamical critical points is therefore more naturally classified by its relaxation-rate spectrum than by its Hamiltonian excitation spectrum.
The TDOS framework introduced here thus elevates relaxation modes to the status of fundamental collective coordinates, providing a principled dynamical basis for infrared universality, memory effects, and emergent macroscopic coherence.
Away from criticality the dynamical spectrum collapses onto a trivial fast mode. Near memory-dominated criticality an extensive continuum of slow collective modes emerges and governs macroscopic coherence. Superconducting pairing in this regime corresponds to condensation within a slow-mode reservoir rather than binding mediated by a single sharp bosonic excitation.
This unified picture resolves long-standing puzzles of unconventional superconductivity, including the coexistence of strange-metal behavior with robust pairing and the apparent universality of dome-shaped phase diagrams across disparate materials. Rather than invoking competing orders or material-specific bosonic glue, these phenomena arise naturally from universal infrared spectral structure.
@@
More broadly, the TDOS framework establishes dynamical spectral organization as a fundamental principle of correlated quantum matter. Memory-dominated criticality represents a new universality class beyond traditional dissipative field theories, with implications extending to non-Fermi liquids, glassy dynamics, and self-organized critical systems.
Future directions include quantitative determination of TDOS spectra in microscopic models and experiments, extension to nonequilibrium driven systems, and exploration of engineered platforms in which slow-mode reservoirs can be dynamically controlled. In this light, high-temperature superconductivity appears not as an anomaly requiring special microscopic ingredients, but as a generic consequence of infrared spectral self-organization in strongly correlated matter.
@@
Memory–dominated dynamics as a guiding principle for room-temperature superconductivity. Recent experimental efforts toward room-temperature superconductivity have largely focused on engineering stronger microscopic pairing interactions through pressure, lattice design, or exotic chemical compositions. Within the present framework, however, enhancing the bare pairing scale alone is insufficient. Our results show that the decisive control parameter for high transition temperatures is not the magnitude of an attractive interaction but the infrared dynamical organization of collective modes encoded in the relaxation-rate density of states.
@@
A dense reservoir of near-marginal slow modes fundamentally reorganizes the Cooper channel, converting moderate intrinsic electronic pairing into an algebraically amplified instability with $T_{c}\sim\rho_{0}E_{\rm pair}$. From this perspective, many existing high-pressure hydride systems and correlated materials may fail to reach extreme transition temperatures not because pairing is weak, but because their collective spectra do not self-organize into a flat TDOS regime.
@@
The present theory therefore suggests a new design principle for superconductivity beyond material-specific glue engineering: the deliberate stabilization of dynamical phases supporting extensive slow-mode continua. Controlling electronic structure, frustration, dimensionality, and feedback mechanisms to promote near-marginal collective dynamics may be as crucial as increasing microscopic attraction itself. Room-temperature superconductivity, in this view, corresponds to a dynamical phase characterized by maximal infrared memory rather than by extreme coupling strength.
@@
We have identified dynamical spectral organization as a fundamental principle governing infrared behavior in strongly correlated quantum matter. Formulating collective dynamics in terms of the time–scale density of states reveals that criticality is generically controlled not by local Markovian dissipation but by the accumulation of slow collective relaxation modes.
A finite infrared TDOS defines a memory–dominated universality class characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic dynamical response. Within this regime, intrinsic electronic pairing tendencies are dynamically amplified, converting the marginal instabilities of conventional theories into algebraic superconducting transitions governed by infrared spectral weight.
This mechanism provides a unified explanation of key experimental phenomena in unconventional superconductors, including superconducting domes, Uemura scaling of transition temperatures with superfluid stiffness, and anomalous long–time dynamical correlations in the normal state. These signatures arise generically from infrared spectral reorganization, without invoking specific bosonic glue, competing orders, or fine tuning.
More broadly, our results recast quantum criticality as a constructive dynamical process that enhances collective quantum phenomena through self–organized memory effects. Memory–dominated criticality thus represents a new paradigm beyond traditional dissipative field theories, with implications extending to non-Fermi liquids, glassy dynamics, and engineered quantum platforms.
In this perspective, high–temperature superconductivity emerges not as an exception requiring special microscopic ingredients, but as a natural consequence of infrared spectral self–organization in strongly correlated matter. Future exploration of TDOS-controlled dynamics may enable systematic design of materials and devices that exploit memory–dominated criticality to stabilize and enhance collective quantum states.
@@
Appendix A Appendix A: Gaussian MSRJD integration of a relaxation-mode reservoir and emergence of the self-energy
@@
In this appendix we derive, without omitting algebraic steps, how a continuum of relaxation (Ornstein–Uhlenbeck) modes generates a nonlocal-in-time memory kernel and an associated self-energy $\Sigma(\omega)$ upon Gaussian functional integration in the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) formalism. The derivation is standard in structure (Gaussian path integration), but we present it in a form tailored to our “slow-mode reservoir” interpretation.
@@
(I) Coupled stochastic dynamics: collective variable plus OU reservoir: We consider a collective degree of freedom $x(t)$ coupled linearly to a continuum of reservoir modes $X_{\lambda}(t)$ labeled by their relaxation rates $\lambda>0$. The coupled stochastic dynamics is
$\dot{x}(t)$ $=-r\,x(t)+\int d\lambda\;g(\lambda)\,X_{\lambda}(t)+\eta(t),$ (S1)
$\dot{X}_{\lambda}(t)$ $=-\lambda\,X_{\lambda}(t)+g(\lambda)\,x(t)+\xi_{\lambda}(t).$ (S2)
@@
$\langle\eta(t)\eta(t^{\prime})\rangle$ $=2D_{x}\,\delta(t-t^{\prime}),$ (S3)
@@
No equilibrium fluctuation–dissipation relation is assumed; $D_{\lambda}$ may be arbitrary. Our goal is to integrate out $\{X_{\lambda}\}$ and obtain an exact effective MSRJD action for (x,x~)($\tilde{x}$) containing a memory kernel.
@@
$Z=\int\mathcal{D}x\,\mathcal{D}\tilde{x}\;\prod_{\lambda}\mathcal{D}X_{\lambda}\,\mathcal{D}\tilde{X}_{\lambda}\;e^{-S[x,\tilde{x},\{X_{\lambda},\tilde{X}_{\lambda}\}]},$ (S6)
@@
$S_{x}$ $=\int dt\;\tilde{x}(t)\!\left[\dot{x}(t)+r\,x(t)-\int d\lambda\;g(\lambda)\,X_{\lambda}(t)\right]$
$\quad-\int dt\;D_{x}\,\tilde{x}(t)^{2},$ (S8)
$S_{\lambda}$ $=\int dt\;\tilde{X}_{\lambda}(t)\!\left[\dot{X}_{\lambda}(t)+\lambda\,X_{\lambda}(t)-g(\lambda)\,x(t)\right]$
@@
$S_{x}\supset-\int dt\;\tilde{x}(t)\,g(\lambda)\,X_{\lambda}(t).$ (S10)
@@
$S^{(\lambda)}[X_{\lambda},\tilde{X}_{\lambda};x,\tilde{x}]=\int dt\;\tilde{X}_{\lambda}(t)\big(\dot{X}_{\lambda}(t)+\lambda X_{\lambda}(t)\big)$ (S11)
$\quad-\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t)-\int dt\;g(\lambda)\,\tilde{x}(t)\,X_{\lambda}(t)$
@@
In what follows, $x,\tilde{x}$ are treated as external sources.
(IV) Functional integration over $X_{\lambda}$: emergence of a delta constraint: Define the linear operator
@@
$S^{(\lambda)}$ $=\int dt\;X_{\lambda}(t)\Big(\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)-g(\lambda)\tilde{x}(t)\Big)$ (S18)
$\quad-\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t)-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}.$
@@
$A(t)\equiv\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)-g(\lambda)\tilde{x}(t).$ (S20)
@@
$\quad\exp\!\left[+\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t)-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}\right].$
The Jacobian factor associated with the functional constraint, $|\det\mathcal{L}_{\lambda}^{\dagger}|^{-1}$, is independent of $x,\tilde{x}$ and can therefore be absorbed into the overall normalization of $Z$.
@@
$\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)=g(\lambda)\tilde{x}(t),\qquad\mathcal{L}_{\lambda}^{\dagger}=-\partial_{t}+\lambda,$ (S22)
@@
Because the constraint is written with the adjoint operator $\mathcal{L}_{\lambda}^{\dagger}=-\partial_{t}+\lambda$, its Green function carries advanced support (equivalently, the retarded kernel will reappear after relabelling, ensuring causality of the $x$-sector response).

### v2

@@ Abstract
This relation reproduces the empirical Uemura scaling observed across cuprates and other unconventional superconductors, where $T_{c}$ tracks the zero-temperature superfluid density. In contrast to weak-coupling BCS superconductors—where phase stiffness is large and essentially decoupled from gap formation—correlated systems operate in a regime in which infrared collective dynamics simultaneously controls pairing and coherence.
Within the present framework, this proportionality is not imposed phenomenologically. It follows directly from the TDOS-controlled infrared fixed point, where a single spectral parameter $\rho_{0}$ governs both the Cooper instability and phase stiffness. The observed scaling of $T_{c}$ with superfluid density thus emerges as a natural consequence of memory-dominated critical dynamics.
@@
Such long–time tails can produce 1/f1/f–type noise spectra and anomalously slow relaxation, phenomena widely reported in strange metals and cuprate superconductors. Unlike conventional overdamped criticality, where correlations decay exponentially and dynamics is effectively Markovian, the memory–dominated regime predicts non-Markovian response with persistent temporal structure. Because the same slow–mode reservoir controls both pairing enhancement and dynamical response, superconducting and normal–state anomalies become two manifestations of a single infrared spectral reorganization.
@@
The TDOS controls: (i) the infrared extent of the slow–mode reservoir, (ii) the dynamical amplification of intrinsic pairing tendencies, (iii) the superfluid stiffness and phase coherence, and (iv) the emergence of non-Markovian long–time response. Quantum criticality thus acts not merely as a source of fluctuations but as a dynamical amplifier governed by infrared spectral reorganization that simultaneously governs thermodynamic and dynamical phenomena in strongly correlated superconductors.
This unified picture yields a concrete strategy for experimental tests: systems that exhibit enhanced low–frequency TDOS weight (or equivalently, strong long–time correlations) are expected to display increased pairing susceptibility, elevated $T_{c}$, and reduced phase stiffness consistent with Uemura scaling. The resulting real-frequency spectral function generated by the memory-dominated self-energy, including its universal scaling form and infrared structure, is derived in Appendix F.
@@
The central result of this work is that infrared collective dynamics in strongly correlated quantum matter is organized by the structure of its relaxation-rate spectrum rather than solely by softening of a single order parameter. The time–scale density of states provides a quantitative measure of this organization. When the TDOS remains finite as $\lambda\to 0$, the system enters a memory–dominated regime characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic infrared response.
@@
This viewpoint clarifies the limits of conventional coarse-grained field theories. Away from criticality, the relaxation spectrum is gapped in the infrared, leading to finite memory times and justifying reduction to local-in-time damping and a small number of collective fields. Near correlated critical regimes, however, relaxation rates condense into a continuous spectrum extending toward zero. In this regime, internal modes cannot be integrated out: the full decay spectrum remains dynamically active across time scales. The breakdown of Markovian coarse-graining is therefore not a model-dependent artifact but a universal consequence of dynamical criticality.
This observation highlights the structural difference from Hertz–Millis–type approaches. Conventional quantum critical theories assume overdamped local dynamics of a small set of bosonic modes, with dissipation controlled by Landau damping. In contrast, the present framework allows for spectral condensation of near-marginal relaxation channels. Criticality is reinterpreted not as single-mode softening but as reorganization of the entire decay spectrum. The TDOS thus replaces phenomenological damping coefficients with a microscopic spectral quantity governing infrared universality.
@@
A direct physical consequence of this spectral reorganization is qualitative enhancement of superconducting instabilities. Superconductivity need not rely on finely tuned bosonic mediators. Instead, intrinsic short-range electronic pairing tendencies are dynamically amplified by the slow-mode reservoir. The exponential suppression characteristic of weak-coupling BCS theory is replaced by algebraic scaling controlled by infrared spectral weight. In this sense, superconducting order corresponds to condensation within an extended slow-mode manifold rather than binding via a sharp bosonic excitation.
@@
A single infrared parameter, $\rho_{0}$, governs both thermodynamic and dynamical phenomena. The superconducting dome reflects truncation of long-time memory away from the optimal spectral reorganization point; Uemura scaling follows from the shared spectral origin of pairing and phase coherence; anomalous normal-state dynamics arises from the same non-Markovian kernel. Quantum criticality therefore acts as a dynamical amplifier linking transport, coherence, and pairing within a unified spectral framework.
The relaxation-rate organization emphasized here is not merely a phenomenological assumption. Both Liouvillian evolution in open quantum systems and generalized memory-kernel dynamics naturally decompose into continua of decay modes characterized by complex relaxation eigenvalues. Whenever long-time correlations are present, Laplace-domain representations generically produce continuous relaxation spectra. A dense TDOS therefore reflects a universal mathematical property of dynamical systems with extended temporal correlations, rather than a model-specific construction.
@@
From this perspective, Hamiltonian energy eigenmodes describe conservative reversible dynamics, whereas relaxation spectra encode the irreversible organization governing collective behavior near criticality. Strongly correlated quantum matter near dynamical critical points may thus be more naturally classified by its relaxation-rate spectrum than by its excitation spectrum alone.
@@
More broadly, the TDOS framework elevates dynamical spectral organization to a fundamental organizing principle of correlated quantum matter. Memory-dominated criticality represents a universality class distinct from traditional dissipative field theories, with potential implications for non-Fermi liquids, glassy systems, and self-organized critical dynamics.
@@
Memory–dominated dynamics as a guiding principle for room-temperature superconductivity. Most contemporary efforts toward room-temperature superconductivity focus on strengthening microscopic pairing interactions through pressure, lattice engineering, or exotic chemical composition. Within the present framework, however, enhancing the bare pairing scale alone is insufficient. The decisive control parameter for high transition temperatures is the infrared dynamical organization of collective modes encoded in the relaxation-rate density of states.
@@
A dense reservoir of near-marginal slow modes reorganizes the Cooper channel, converting moderate intrinsic electronic pairing into an algebraically amplified instability with $T_{c}\sim\rho_{0}E_{\rm pair}$. From this viewpoint, certain high-pressure hydrides or correlated materials may fail to reach extreme transition temperatures not because pairing is weak, but because their collective spectra do not self-organize into a flat TDOS regime.
@@
The present theory therefore suggests a complementary design principle for superconductivity beyond material-specific glue engineering: stabilization of dynamical phases that support extensive slow-mode continua [27]. Control of electronic frustration, dimensionality, feedback mechanisms, or proximity to dynamical criticality may be as crucial as increasing microscopic attraction itself. Room-temperature superconductivity, in this perspective, corresponds to a dynamical phase characterized by maximal infrared memory rather than simply by extreme coupling strength.
@@
We have identified dynamical spectral organization as a fundamental principle governing infrared behavior in strongly correlated quantum matter. By formulating collective dynamics in terms of the time–scale density of states, we show that criticality is generically controlled not by local Markovian dissipation but by the accumulation of slow collective relaxation modes.
A finite infrared TDOS defines a memory–dominated universality class characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic dynamical response. Within this regime, intrinsic electronic pairing tendencies are dynamically amplified, converting marginal logarithmic instabilities into algebraic superconducting transitions governed by infrared spectral weight.
This mechanism provides a unified account of key experimental phenomena in unconventional superconductors, including superconducting domes, Uemura scaling of transition temperatures with superfluid stiffness, and anomalous long–time dynamical correlations in the normal state. These signatures arise naturally from infrared spectral reorganization, without requiring specific bosonic mediators or material-dependent fine tuning.
More broadly, the results recast quantum criticality as a dynamical process that enhances collective quantum phenomena through extended temporal correlations. Memory–dominated criticality therefore represents a universality class distinct from traditional dissipative field theories, with implications for non-Fermi liquids, glassy dynamics, and engineered quantum platforms.
In this perspective, high–temperature superconductivity appears as a generic consequence of infrared spectral organization in strongly correlated matter. Systematic exploration of TDOS-controlled dynamics may provide a pathway toward stabilizing and enhancing collective quantum states in future materials and devices.
@@
Appendix A Appendix A: Relaxation-rate spectra as the fundamental dynamical basis of collective dynamics
In conventional quantum many-body theory, collective behavior is often organized in terms of energy eigenmodes of an underlying Hamiltonian,
$H|n\rangle=E_{n}|n\rangle,$ (S1)
with dynamical response expressed through oscillatory Fourier components at frequencies set by energy differences. This viewpoint is natural for closed systems whose dynamics is unitary and time-reversal invariant, and it emphasizes stationary states and conservative motion.
However, energy eigenmodes provide only a special organizing basis. In generic interacting settings—especially in open, driven, or internally coarse-grained systems—the physically relevant long-time organization is not set by Hamiltonian energies but by the stability spectrum of the effective dynamical generator. In this work, the central spectral object is therefore the distribution of relaxation rates governing irreversible decay and memory.
(I) Dynamical linearization and relaxation modes: For a general high-dimensional (possibly coarse-grained) dynamical system,
$\dot{x}(t)=F(x(t)),$ (S2)
linearization about a metastable configuration $x^{*}$ yields
The eigenvalue problem
$Av_{i}=\lambda_{i}v_{i}$ (S4)
$\delta x_{i}(t)\sim e^{-\lambda_{i}t}.$ (S5)
The real parts of $\lambda_{i}$ set decay time scales, while imaginary parts encode oscillatory components. Purely oscillatory energy-mode dynamics corresponds to the singular limit $\mathrm{Re}\,\lambda_{i}=0$ for all modes, which is realized only in strictly Hamiltonian, non-dissipative evolution.
Thus the relaxation-rate spectrum $\{\lambda_{i}\}$ constitutes the natural dynamical basis for generic collective phenomena, including metastability, critical slowing down, and non-Markovian memory.
(II) Liouvillian formulation (closed, open, and non-Markovian dynamics): A conceptually complete formulation of nonequilibrium quantum dynamics is given in terms of the density matrix $\hat{\varrho}(t)$ and its dynamical generator. For a closed system, $\hat{\varrho}(t)$ evolves by the von Neumann equation,
$\frac{d\hat{\varrho}}{dt}=-i[H,\hat{\varrho}]\equiv\mathcal{L}_{H}\hat{\varrho}$, (S6)
where $\mathcal{L}_{H}$ is the Hamiltonian Liouvillian superoperator. If $H|n\rangle=E_{n}|n\rangle$, then the operator basis $|n\rangle\langle m|$ diagonalizes $\mathcal{L}_{H}$:
$\mathcal{L}_{H}\big(|n\rangle\langle m|\big)=-i(E_{n}-E_{m})\,|n\rangle\langle m|,$ (S7)
so the Liouvillian eigenvalues are
$\lambda_{nm}=-i(E_{n}-E_{m}),$ (S8)
lying purely on the imaginary axis and reproducing undamped oscillations.
For open quantum systems, the reduced density matrix typically obeys a master equation with a dissipative contribution. In the Markovian limit, the evolution takes Lindblad form,
[…]
where the jump operators $L_{\mu}$ encode decoherence and dissipation. The resulting Liouvillian $\mathcal{L}$ is generically non-Hermitian as a superoperator and therefore possesses complex eigenvalues
$\mathcal{L}R_{\alpha}=\lambda_{\alpha}R_{\alpha},\qquad\mathrm{Re}\,\lambda_{\alpha}\leq 0,$ (S10)
with decay rates set by $\mathrm{Re}\,\lambda_{\alpha}$. The density matrix admits the modal expansion
$\hat{\varrho}(t)=\hat{\varrho}_{\rm ss}+\sum_{\alpha\neq 0}c_{\alpha}\,e^{\lambda_{\alpha}t}\,R_{\alpha},$ (S11)
where $\hat{\varrho}_{\rm ss}$ is the stationary state ($\lambda_{0}=0$) and the long-time approach is controlled by the Liouvillian spectral gap and any near-marginal continuum.
More generally, when the environment or eliminated internal coordinates retain memory, the reduced dynamics becomes non-Markovian and takes a memory-kernel form (Nakajima–Zwanzig structure),
$\frac{d\hat{\varrho}}{dt}=\int_{0}^{t}ds\;\mathcal{K}(t-s)\,\hat{\varrho}(s),$ (S12)
where $\mathcal{K}(t)$ is a causal superoperator kernel. In Laplace space,
$\tilde{\hat{\varrho}}(s)=\big[s-\tilde{\mathcal{K}}(s)\big]^{-1}\hat{\varrho}(0),$ (S13)
so the poles and branch cuts of $s-\tilde{\mathcal{K}}(s)$ define the intrinsic relaxation spectrum. In this sense, even for quantum systems the relevant organizing structure for long-time dynamics is generically a relaxation-rate spectrum (possibly continuous), not an energy spectrum.
(III) Memory kernels and continuous relaxation spectra: The same spectral logic appears already at the level of linear causal dynamics for a collective coordinate $\phi(t)$:
$\dot{\phi}(t)+\int_{0}^{t}ds\,K(t-s)\phi(s)=\eta(t).$ (S14)
Taking the Laplace transform gives
$\tilde{\phi}(s)=\frac{\tilde{\eta}(s)}{s+\tilde{K}(s)}.$ (S15)
The analytic structure of $s+\tilde{K}(s)$ determines the intrinsic relaxation content: isolated poles correspond to discrete decay modes, while branch cuts encode continua of decay rates. Whenever long-time memory is present, the causal kernel admits a spectral representation of the form
$K(t)=\int_{0}^{\infty}d\lambda\,\rho(\lambda)e^{-\lambda t},$ (S16)
which defines the time-scale spectrum. A flat infrared TDOS produces the universal tail $K(t)\sim 1/t$.
Thus continuous relaxation spectra arise generically from causality and memory, independent of microscopic Hamiltonian structure.
(IV) Energy modes as a special limit: The familiar Fourier-mode picture of conservative dynamics corresponds to the singular limit in which the relevant spectrum collapses onto the imaginary axis. In the closed-system Liouvillian case this appears as $\lambda_{nm}=-i(E_{n}-E_{m})$, while in the memory-kernel representation it corresponds to eliminating all decay rates (no weight at $\mathrm{Re}\,\lambda>0$), recovering purely oscillatory behavior. Away from this limit, generic interacting and/or open systems naturally develop broad spectra of relaxation rates controlling infrared dynamics.
(V) Physical implication: The TDOS—the infrared distribution of collective relaxation rates— is therefore the fundamental organizing variable of collective dynamics in correlated systems. Energy eigenmodes describe an integrable, conservative limit, whereas relaxation-rate spectra capture the universal structure of irreversible, critical, and strongly interacting matter.
Memory-dominated criticality corresponds precisely to the formation of an extensive continuum of near-marginal relaxation modes. The resulting TDOS controls dynamical response, pairing enhancement, and infrared universality, providing a deeper and more generally applicable foundation than energy-based descriptions.
@@ Appendix A Appendix A: Relaxation-rate spectra as the fundamental dynamical basis of collective dynamics
Appendix B Appendix B: Gaussian MSRJD integration of a relaxation-mode reservoir and emergence of the self-energy
@@
In this appendix we derive how a continuum of relaxation (Ornstein–Uhlenbeck) modes generates a nonlocal-in-time memory kernel and an associated self-energy $\Sigma(\omega)$ upon Gaussian functional integration in the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) formalism. The derivation is standard in structure (Gaussian path integration), but we present it in a form tailored to our “slow-mode reservoir” interpretation.
@@
(I) Coupled stochastic dynamics: We consider a collective degree of freedom $\phi(t)$ coupled linearly to a continuum of reservoir modes $X_{\lambda}(t)$ labeled by their relaxation rates $\lambda>0$. The coupled stochastic dynamics is
$\dot{\phi}(t)$ $=-r\,\phi(t)+\int d\lambda\;g(\lambda)\,X_{\lambda}(t)+\eta(t),$ (S17)
$\dot{X}_{\lambda}(t)$ $=-\lambda\,X_{\lambda}(t)+g(\lambda)\,\phi(t)+\xi_{\lambda}(t).$ (S18)
@@
$\langle\eta(t)\eta(t^{\prime})\rangle$ $=2D_{\phi}\,\delta(t-t^{\prime}),$ (S19)
@@
[…]
@@
$Z=\int\mathcal{D}\phi\,\mathcal{D}\tilde{\phi}\;\prod_{\lambda}\mathcal{D}X_{\lambda}\,\mathcal{D}\tilde{X}_{\lambda}\;e^{-S[\phi,\tilde{\phi},\{X_{\lambda},\tilde{X}_{\lambda}\}]},$ (S22)
@@
$S_{\phi}$ $=\int dt\;\tilde{\phi}(t)\!\left[\dot{\phi}(t)+r\,\phi(t)-\int d\lambda\;g(\lambda)\,X_{\lambda}(t)\right]$
$\quad-\int dt\;D_{\phi}\,\tilde{\phi}(t)^{2},$ (S24)
$S_{\lambda}$ $=\int dt\;\tilde{X}_{\lambda}(t)\!\left[\dot{X}_{\lambda}(t)+\lambda\,X_{\lambda}(t)-g(\lambda)\,\phi(t)\right]$
@@
$S_{\phi}\supset-\int dt\;\tilde{\phi}(t)\,g(\lambda)\,X_{\lambda}(t).$ (S26)
@@
$S^{(\lambda)}[X_{\lambda},\tilde{X}_{\lambda};\phi,\tilde{\phi}]=\int dt\;\tilde{X}_{\lambda}(t)\big(\dot{X}_{\lambda}(t)+\lambda X_{\lambda}(t)\big)$ (S27)
$\quad-\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t)-\int dt\;g(\lambda)\,\tilde{\phi}(t)\,X_{\lambda}(t)$
@@
In what follows, $\phi,\tilde{\phi}$ are treated as external sources.
(IV) Functional integration over $X_{\lambda}$: Define the linear operator
@@
$S^{(\lambda)}$ $=\int dt\;X_{\lambda}(t)\Big(\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)-g(\lambda)\tilde{\phi}(t)\Big)$ (S34)
$\quad-\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t)-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}.$
@@
$A(t)\equiv\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)-g(\lambda)\tilde{\phi}(t).$ (S36)
@@
$\quad\exp\!\left[+\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t)-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}\right].$
The Jacobian factor associated with the functional constraint, $|\det\mathcal{L}_{\lambda}^{\dagger}|^{-1}$, is independent of $\phi,\tilde{\phi}$ and can therefore be absorbed into the overall normalization of $Z$.
@@
$\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)=g(\lambda)\tilde{\phi}(t),\qquad\mathcal{L}_{\lambda}^{\dagger}=-\partial_{t}+\lambda,$ (S38)
@@
Because the constraint is written with the adjoint operator $\mathcal{L}_{\lambda}^{\dagger}=-\partial_{t}+\lambda$, its Green function carries advanced support (equivalently, the retarded kernel will reappear after relabelling, ensuring causality of the $\phi$-sector response).